# A Weighted Loss Approach to Robust Federated Learning under Data Heterogeneity

### v2

@@ B.1 Robust Distributed Stochastic Heavy Ball Algorithm
We present here the robust distributed stochastic Heavy Ball algorithm. More details can be found in [3]. We use superscript in brackets to denote the iteration index, i.e., (t), and subscripts to indicate the worker index.
Input: Initial model $\theta^{(0)}$, initial momentum $m_{i}^{(0)}=0$ for honest workers, robust aggregation $F$, learning rate $\gamma$, momentum coefficient $\beta$, and number of steps $T$.
for $t=1,\ldots,T$ do
       Server broadcasts $\theta^{(t-1)}$ to all workers;
       for every honest worker $i$ do
             Compute a stochastic gradient $g^{(t)}_{i}$;
             Update local momentum $m_{i}^{(t)}\leftarrow\beta m_{i}^{(t-1)}+(1-\beta)g^{(t)}_{i}$;
             Send $m_{i}^{(t)}$ to the server;
       end for
      Server aggregates the momentums:
           $R^{(t)}=F(m_{1}^{(t)},\ldots,m_{n}^{(t)})$;
       Server updates the model: $\theta^{(t)}=\theta^{(t-1)}-\gamma R^{(t)}$;
end for
@@
B.2 Model Architectures
We adopt the following abreviations : L(#outputs) denotes a fully connected linear layer; R is a ReLU activation; T is a Tanh activation; C(#channels) represents a 2D convolutional layer with kernel size $5$, padding 00, and stride $1$; M denotes 2D max-pooling with kernel size $2$; B stands for batch normalization; and D represents dropout with a fixed probability of $0.25$.
@@ B.2 Model Architectures
  • MNIST and Fashion MNIST (CNN from [3]): C(20)-R-M–C(20)-R-M–L(500)-R–L(10)
  • CIFAR10 (CNN from [3]):(3,32×32)-C(64)-R-B-C(64)-R-B-M-D-C(128)-R-B-C(128)-R-B-M-D-L(128)-R-D-L(10)
  • Purchase100 (fully connected neural networks from [31]): L(1024)-L(512)-T-L(256)-T-L(100)
@@
  • MNIST and Fashion MNIST (from [3]): $\gamma_{t}=0.75/(1+\left\lfloor\frac{t}{50}\right\rfloor)$
  • CIFAR10 (from [3]) and Purchase100: $\gamma_{t}=0.25$ if $t\leq 1500$, $0.025$ otherwise.
B.3 Mimic with Most Surrounded Outlier Selection
@@
where $\operatorname{dist}$ denotes the Euclidean distance. For any vector $\mathbf{v}_{j}$, there exists an index $r$ such that $\sigma(i,r)=j$, meaning that $\mathbf{v}_{j}$ is the $r$‑th farthest vector from $\mathbf{v}_{i}$. Note that $\sigma(i,h)=i$. We define the operator $\operatorname{rank}$ as:
@@
Input: Honest updates $\mathbf{v}_{1},\ldots,\mathbf{v}_{h}\in\mathbb{R}^{d}$, number of Byzantine workers $f$
Output: Update $\mathbf{v}_{\text{argmax}}$ with highest score
for $i=1,\ldots,h$ do
       $s_{i}\leftarrow 0$ ;
end for
$\textit{center}\leftarrow\frac{1}{h}\sum_{i=1}^{h}\mathbf{v}_{i}$ ;
for $i=1,\ldots,h$ do
       for $j=1,\ldots,h$ do
             $r\leftarrow\operatorname{rank}(\mathbf{v}_{i},\mathbf{v}_{j})$ ;
             $s_{j}\leftarrow s_{j}+\min(f,r)\cdot\operatorname{dist}(\mathbf{v}_{j},\textit{center})$ ;
       end for
end for
$k\leftarrow\arg\max_{j}s_{j}$ return $\mathbf{v}_{k}$
@@
B.4 Comprehensive Results
We report the full experimental results. For MNIST, see Tables 5, 6, and 7; for Fashion MNIST, see Tables 8, 9, and 10; for Purchase100, see Tables 11, 12, and 13; and for CIFAR10, see Tables 14, 15, and 16. Each set of three tables corresponds to low, medium, and high heterogeneity levels ($\alpha=3$, $1$, and $0.3$).

### v3

@@ Appendix G Experimental Details
We present here the robust distributed stochastic Heavy Ball algorithm. More details can be found in (Allouah et al. 2023). We use superscript in brackets to denote the iteration index, i.e., (t), and subscripts to indicate the worker index.
@@
0: Initial model $\theta^{(0)}$, initial momentum $m_{i}^{(0)}=0$ for honest workers, robust aggregation $F$, learning rate $\gamma$, momentum coefficient $\beta$, and number of steps $T$
1:for $t=1$ to $T$ do
2:  Server broadcasts $\theta^{(t-1)}$ to all workers
3:  for every honest worker $i$ do
4:   Compute a stochastic gradient $g^{(t)}_{i}$
5:   Update local momentum: $m_{i}^{(t)}\leftarrow\beta m_{i}^{(t-1)}+(1-\beta)g_{i}^{(t)}$
6:   Send $m_{i}^{(t)}$ to the server
7:  end for
8:  Server aggregates the momenta:
9:    $R^{(t)}=F(m_{1}^{(t)},\ldots,m_{n}^{(t)})$
10:  Server updates the model:
11:    $\theta^{(t)}=\theta^{(t-1)}-\gamma R^{(t)}$
12:end for
Model Architectures
We adopt the following abreviations : L(#outputs) denotes a fully connected linear layer; R is a ReLU activation; T is a Tanh activation; C(#channels) represents a 2D convolutional layer with kernel size $5$, padding 0, and stride $1$; M denotes 2D max-pooling with kernel size $2$; B stands for batch normalization; and D represents dropout with a fixed probability of $0.25$.
@@
  • MNIST and Fashion MNIST (CNN from (Allouah et al. 2023)): C(20)-R-M–C(20)-R-M–L(500)-R–L(10)
  • CIFAR10 (CNN from (Allouah et al. 2023)):(3,32×32)-C(64)-R-B-C(64)-R-B-M-D-C(128)-R-B-C(128)-R-B-M-D-L(128)-R-D-L(10)
  • Purchase100 (fully connected neural networks from (Nasr, Shokri, and Houmansadr 2018)): L(1024)-L(512)-T-L(256)-T-L(100)
@@
  • MNIST and Fashion MNIST (from (Allouah et al. 2023)): $\gamma_{t}=0.75/(1+\left\lfloor\frac{t}{50}\right\rfloor)$
  • CIFAR10 (from (Allouah et al. 2023)) and Purchase100: $\gamma_{t}=0.25$ if $t\leq 1500$, $0.025$ otherwise.
Mimic with Most Surrounded Outlier Selection
@@
where $\operatorname{dist}$ denotes the Euclidean distance. For any vector $\mathbf{v}_{j}$, there exists an index $r$ such that $\sigma(i,r)=j$, meaning that $\mathbf{v}_{j}$ is the rr‑th farthest vector from $\mathbf{v}_{i}$. Note that $\sigma(i,h)=i$. We define the operator $\operatorname{rank}$ as:
@@
0: Honest updates $\mathbf{v}_{1},\ldots,\mathbf{v}_{h}\in\mathbb{R}^{d}$, number of Byzantine workers $f$
0: Update $\mathbf{v}_{\text{argmax}}$ with highest score
1:for $i=1$ to $h$ do
2:  $s_{i}\leftarrow 0$
3:end for
[…]
5:for $i=1$ to $h$ do
6:  for $j=1$ to $h$ do
7:   $r\leftarrow\operatorname{rank}(\mathbf{v}_{i},\mathbf{v}_{j})$
8:   $s_{j}\leftarrow s_{j}+\min(f,r)\cdot\operatorname{dist}(\mathbf{v}_{j},\textit{center})$
9:  end for
10:end for
[…]
12:return  $\mathbf{v}_{k}$
Appendix H Full and Additional Results
We report summarized performance in Table 4 for MNIST and Purchase100, which show similar method behaviors as Table 1. Additionally, full experimental results are provided. For MNIST, see Tables 56, and 7; for Fashion MNIST, see Tables 89, and 10; for Purchase100, see Tables 1112, and 13; and for CIFAR10, see Tables 1415, and 16. Each set of three tables corresponds to low, medium, and high heterogeneity levels ($\alpha=3$, $1$, and $0.3$).
Table 4: Test accuracy (± standard deviation) in %, averaged over training, on MNIST and Purchase100. Global results are averaged across aggregators (CWMed, CwTM, GM, MKrum) and attacks (ALIE, FOE, LF, Mimic, SF). Worst-case results correspond to the worst attack, averaged across aggregators. Colors rank values within columns: yellow (1st), blue (2nd), red (3rd). See full table in Appendix.
(a) MNIST (Global)
Accuracy
Accuracy Gain Over Top Non-WoLA Methods
WoLA 0.0 0.0 0.0 +6.4 0.0 +0.4 +2.4 +17.3 -0.4 +1.5 +21.4 +16.5
WoLA+NNM 0.0 +0.2 +0.6 +8.8 +0.1 +0.6 +3.1 +19.9 -0.2 +1.8 +22.5 +26.6
WoLA 0.0 0.0 -0.2 +5.9 -0.1 +0.3 +2.0 +13.6 -0.4 +1.4 +20.5 +9.3
WoLA+NNM 0.0 +0.1 +0.3 +8.3 -0.1 +0.5 +2.8 +10.3 -0.2 +1.8 +22.3 +10.7
(b) Purchase100 (Global)
Accuracy
Accuracy Gain Over Top Non-WoLA Methods
WoLA -1.0 +1.3 +5.4 +14.6 -1.4 +3.1 +9.3 +14.8 -6.4 -1.7 +14.6 +6.8
WoLA -1.2 +0.8 +4.0 +10.2 -1.6 +2.4 +7.2 +5.2 -6.7 -2.9 +11.9 -3.9
(c) MNIST (Worst-case)
Accuracy
Accuracy Gain Over Top Non-WoLA Methods
WoLA -0.1 +0.1 -0.8 +19.7 -0.1 +1.3 +6.0 +41.9 -0.4 +6.5 +28.1 +10.7
WoLA+NNM 0.0 +0.3 +1.3 +23.5 0.0 +1.6 +8.3 +44.3 -0.1 +7.0 +29.3 +6.0
WoLA -0.1 0.0 -0.8 +20.3 -0.1 +1.3 +5.7 +41.3 -0.4 +6.4 +26.3 +6.5
WoLA+NNM +0.1 +0.2 +0.9 +23.2 -0.1 +1.5 +7.9 +28.9 -0.1 +7.0 +30.6 -3.2
(d) Purchase100 (Worst-case)
Accuracy
Accuracy Gain Over Top Non-WoLA Methods
WoLA -0.7 +3.4 +9.2 +29.0 +0.8 +10.0 +15.4 +16.1 -1.5 +10.4 +15.1 +3.6
WoLA -0.9 +2.8 +7.6 +23.2 +0.5 +9.1 +12.1 +18.4 -1.8 +8.9 +20.3 +1.7